OpenStudy (anonymous):

Find the partial fraction decomposition: (4x-11)/((2x^2)-x-3)

sam (.sam.):

(4x-11)/((2x^(2))-x-3) (4x-11)/(2x^(2)-x-3) (4x-11)/((x+1)(2x-3)) --------------------------------- (A)/(x+1)+(B)/(2x-3) 4=2A+B -11=-3A+1B -------------------------------------- 4=2A+B -11=-3A+1B Simultaneous B=-2 A=3 (3)/(x+1)-(2)/(2x-3)

OpenStudy (anonymous):

thanks!